mark is 10 yrs younget than larry.larrys age8 years from now will wxceed twice marks age3 yearsago by 4 years . how old is each boy?
Answer by ankor@dixie-net.com(22740)

mark is 10 yrs younger than larry.
M = L - 10
:
larry's age 8 years from now, will exceed twice marks age 3 years ago by 4 years.
L + 8 = 2(M-3) + 4
Do some basic algebra on this
L + 8 = 2M - 6 + 4
L + 8 = 2M - 2
L = 2M - 2 - 8
L = 2M - 10
:
how old is each boy?
Substitute (L-10) for M in the above equation, solve for L:
L = 2(L-10) - 10
L = 2L - 20 - 10
L = 2L - 30
L - 2L = -30
-L = -30
L = -30/-1
L = +30 is Larry's age
Then
M = 30 - 10 = 20 is Marks age
:
:
Check our solution using the statememt:
"larrys age8 years from now will exceed twice marks age 3 years ago by 4 years".
(30 + 8) = 2(20-3) + 4
38 = 2(17) + 4; confirms our solution
:
Did this method makes sense to you? Any questions?